Fermi National Laboratory

Accelerator Update for January 28, 2008 - January 30, 2008


Monday January 28

The day shift began with Operations monitoring the Tevatron (TeV) store (92.27E30) and the Recycler stash (198.3E10), with the Antiproton Source stacking (32.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.

Water technicians reported at 9:40 AM that they were beginning work on air compressors in the following TeV sectors: B4, C1, C2, and C3. This work took about four hours.

Fire techs reported at 9:42 AM that they were testing the D0 fire panels. These tests took about 15 minutes. They continued these tests around the TeV ring throughout the day.

At 10:53 AM, Operations reported a problem with beam positions in the TeV. There was no tuning and no indication that there was an earthquake somewhere, but there was a dip in luminosity. Operators contacted a TeV expert.

At 12:32 PM, a Linac RF station (LRF2) tripped off. It reset.

At 12:41 PM, a MI expert began a collimator study.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons five times during the day shift.

At 4:28 PM, a Linac Klystron RF station (KRF2) tripped off. It reset.

At 6:53 PM, KRF6 tripped off. It reset.

At 7:50 PM, KRF7 tripped off. It reset.

At 9:15 PM, KRF4 tripped off. It reset.

Operations terminated store 5871 at 10:09 PM.

Operations put the TeV into shot setup at 10:41 PM.

At 11:23 PM, Operations reported problems with the TeV T117 orbit application that uses BPM data. Operations contacted experts.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Tuesday January 29

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (432.7E10), with Pbar stacking (93.7E10), and with NuMI and BNB taking beam.

A TeV expert helped operators restore the T117 BPM data.

At 1:33 AM, a MI Lambertson (I:LAM62) tripped off.

At 1:46 AM, Operations reported that a TeV collimator at D17 did not go to its nominal position. Operations had to move the collimators out and then back in to their nominal positions.

Operations established store 5872 at 2:03 AM with an initial luminosity of 266.35E30.

At 3:15 AM, LRF4 tripped off. It reset.

At 6:36 AM, BNB lost its beam permit for the first of many times due to a low current reading (on E:MBEXIP).

At 7:02 AM, a TeV corrector magnet power supply (T:VE16) tripped off. The store remained intact.

Operations reported at 7:56 AM that the Recycler had stashed five times during the midnight shift.

Water techs reported at 8:11 AM that they would be working on the TeV air compressors in the following sectors during the morning: C4, E1, and E2.

Operations reported at 8:47 AM that Switchyard (SY) experts were beginning to run beam to the SY absorber.

At 9:30 AM, KRF7 tripped off. It reset.

Pbar experts began a Pbar Debuncher study at 10:35 AM.

At 10:46 AM, Operations reported that the MI collimator studies caused the Pelletron to trip off. Operators contacted experts who said to leave the Electron Cooling off.

Operations reported at 10:55 AM that Pbar had actually surpassed their stacking record with a rate of 25.19mA for an hour.

The MI collimator studies ended and SY experts established beam to the SY absorber at 11:27 AM.

At 11:44 AM, KRF7 tripped off for the first of two times in five minutes. It reset.

At 12:12 PM, an E:MBEXIP current peak trip caused BNB to trip off for the first of eight times in about 90 minutes. Operators contacted an expert.

At 1:36 PM, store 5872 quenched at sector D0. Operators investigated.

At 1:41 PM, BRF16 tripped off.

Booster experts began tuning their accelerator at 1:47 PM.

Operations reported at 1:48 PM that a quadrupole power supply (T:D0Q6) tripped off before the quench on a ground fault. Experts began a hunt for a TeV ground fault.

Power supply experts reported at 2:24 PM that they had changed the regulator for T:VE16.

At 2:52 PM, Operations reported that NuMI tripped off due to a number of power supplies tripping off from an LCW problem. Operators contacted experts who reported that the trip was due to a low LCW expansion tank at MI-60. Water Techs will fill the tank. NuMI was back online by 3:25 PM.

Operations reported at 4 PM that the Recycler had stashed four times during the day shift.

Operations prepared D0 for an access at 4:01 PM.

Pbar experts reported at 4:44 PM that they had resolved their Debuncher cooling problems.

D0 came out of access at 5:04 PM.

At 7:23 PM, the TeV sector D4 interlocks dropped. Operators investigated and found that a D4 emergency exit door was loose and was probably rattled by the wind, which caused the interlocks to drop. Operators searched and secured the area.

Operations reported at 10:51 PM that experts discovered the LCW supply valve for the power leads at C47 was closed. When the valve was opened, the fault improved, but didn’t clear. This means that the magnet probably did not have a ground fault.

Wednesday January 30

The midnight shift began with a TeV hipot and turn on occurring, with Operations monitoring the stash (341.7E10), with Pbar stacking (162.8E10), and with NuMI and BNB taking beam.

At 12:48 AM, Operations reported that Booster beam positions went bad. Operators investigated.

Operations reported at 1:13 AM that the TeV hipot was good.

At 1:30 AM, Operations reported having problems with Booster injection.

At 1:57 AM, Operations reported that the 400 MeV area beam permit had tripped off due to bad beam positions. Operators discovered that the 400 MeV chopper power supply faulted off due to a termination problem. Operators called in experts.

Operations put Pbar into standby at 2:06 AM.

Experts reported at 3:20 AM that the 400 MeV area chopper cable was too short to reterminate; they would have to use one of the spare cables.

Operations prepared Linac for an access at 3:21 AM to repair the chopper. The access ended at 4:48 AM.

Chopper repairs ended and Linac resumed sending beam at 5:23 AM.

Operations resumed sending beam to NuMI, BNB, and Pbar by 5:25 AM.

At 5:29, LRF2 tripped off. It reset.

Operations put the TeV into shot setup at 5:34 AM.

Operations turned off beam to NuMI at 5:57 AM for a scheduled access.

At 5:58 AM, LRF2 tripped off for the first of four times in 23 minutes. It reset.

At 5:59 AM, Operations reported that the TeV BPMs changed modes again. Operators rebooted and restored the processor for collider mode. This delayed shot setup.

The Plans for Wednesday

The plans for today are to establish a TeV store and then go into a 10-12 hour shutdown to replace the bad MI quadrupole magnet D631. Booster and Linac experts will access and conduct maintenance work.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/30/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory